\begin{thebibliography}{99}


\bibitem{besse} Besse, A. L. \textit{Einstein manifolds}. Springer Science and Business Media, 2007.

\bibitem{bishop} Bishop, R. L.; O'Neill, B. \textit{Manifolds of negative curvature}. Transactions of the American Mathematical Society, v. 145, p. 1-49, 1969.


\bibitem{brovaz} Brozos-Vázquez, M.; García-Río, E.; Vázquez-Lorenzo, R. \textit{Warped product metrics and locally conformally flat structures.} Matemática Contemporânea. 28(5):91-110, 2005.



\bibitem{cachen} Cao, H. D.; Chen, Q. \textit{On locally conformally flat gradient steady Ricci solitons}. Transactions of the American Mathematical Society, 364(5), pp.2377-2391, 2012.

\bibitem{caoqian} Cao, H. D.; Chen, Q. \textit{On Bach-flat gradient shrinking Ricci solitons}. Duke Mathematical Journal 162, no. 6, 1149-1169, 2013.

\bibitem{cccmm} Cao, H. D.; Catino, G.; Chen, Q.; Mantegazza, C.; Mazzieri, L. \textit{Bach-flat gradient steady Ricci solitons}. Calculus of Variations and Partial Differential Equations, 49, pp.125-138, 2014.

\bibitem{caozhou} Cao, H. D.; Zhou, D. \textit{On complete gradient shrinking Ricci solitons}. Journal of Differential Geometry, v. 85, n. 2, p. 175-186, 2010.


\bibitem{choui} Chouikha, A. R. \textit{Existence of Metrics with Harmonic Curvature and Non Parallel Ricci Tensor}. Balkan Journal of Geometry and Its Applications 8, no. 2, 21-30, 2003.

\bibitem{ivey2} Ivey, T.: \textit{Local existence of Ricci solitons}. Manuscripta Math. 91, 151–162, 1996.

\bibitem{danwan} Dancer, A. S.; Wang M. Y.; \textit{Some new examples of non-kähler ricci solitons.} Mathematical Research Letters 16, no. 2, 2009.

\bibitem{derd} Derdziński, A. \textit{Classification of certain compact Riemannian manifolds with harmonic curvature and non-parallel Ricci tensor}. Mathematische Zeitschrift, 172(3), 273-280, 1980.


\bibitem{DobUn} Dobarro, F.; Ünal, B. \textit{Curvature of multiply warped products}. Journal of Geometry and Physics, 55(1), pp.75-106, 2005.

\bibitem{fega} Fernández-López, M; García-Río, E. \textit{Rigidity of shrinking Ricci solitons}. Mathematische Zeitschrift, 269, pp.461-466, 2011.



\bibitem{kim1} Kim, J. \textit{On a classification of 4-d gradient Ricci solitons with harmonic Weyl curvature}. The Journal of Geometric Analysis, 27(2), 986-1012, 2017.

\bibitem{kim3} Kim, J. \textit{Classification of Gradient Ricci Solitons with Harmonic Weyl Curvature}. The Journal of Geometric Analysis, 35(5), p.139, 2025.

\bibitem{feng} Li, F. \textit{Rigidity of complete gradient steady Ricci solitons with harmonic Weyl curvature}. Pacific Journal of Mathematics, 335(2), pp.323-353, 2025.



\bibitem{muse} Munteanu, O.; Sesum, N. \textit{On gradient Ricci solitons}. Journal of geometric analysis, 23, pp.539-561, 2013.

\bibitem{oneill} O'Neill, B. \textit{Semi-Riemannian Geometry With Applications to Relativity} 103. Academic press, 1983.

\bibitem{shin} Shin, J. \textit{On the classification of 4-dimensional $(m,\rho)$-quasi-Einstein manifolds with harmonic Weyl curvature}. Annals of Global Analysis and Geometry, 51(4), 379-399, 2017.

\bibitem{catinoBook} Catino, G.; Mastrolia, P. \textit{A Perspective on Canonical Riemannian Metrics}. Progress in Mathematics, 336. Birkhäuser, 2020.


\end{thebibliography}

\documentclass[reqno]{amsproc}
\usepackage[utf8]{inputenc}
\usepackage{orcidlink}
\usepackage{amsfonts}
\usepackage{graphicx}
\usepackage{verbatim}
\usepackage{amscd}
\usepackage{amsmath, physics}
\usepackage{enumitem}
\usepackage{amssymb}
\usepackage{latexsym}
\usepackage{hyperref}
\usepackage[all]{xy}
\usepackage{color}
\usepackage{mathrsfs}
\newtheorem{theorem}{\bf Theorem}
\newtheorem{proposition}{\bf Proposition}
\newtheorem{corollary}{\bf Corollary}
\newtheorem{lemma}{\bf Lemma}
\theoremstyle{definition}
\newtheorem{remark}{\bf Remark}

\usepackage{empheq,etoolbox}
\usepackage{stmaryrd}
\numberwithin{equation}{section}
\patchcmd{\subequations}% <cmd>
  {\theparentequation\alph{equation}}% <search>
  {\theparentequation.\alph{equation}}% <replace>
  {}{}% <success><failure>

\begin{document}

\title[The Ricci tensor of a gradient Ricci soliton with harmonic Weyl tensor]{The Ricci tensor of a gradient Ricci soliton with harmonic Weyl tensor}

\author[V. Borges]{V. Borges$^1$}
\author[M. A. R. M. Horácio]{M. A. R. M. Horácio$^2$}
\author[J. P. dos Santos]{J. P. dos Santos$^3$}

\address{$^1$Institute of Exact and Natural Sciences, Universidade Federal do Pará, Belém, Pará, Brazil}

\address{$^{2,3}$Mathematics Department, Universidade de Brasília, Brasília, Distrito Federal, Brazil}

\email{$^1$valterborges@ufpa.br}
\email{$^2$m.a.r.m.horacio@mat.unb.br}
\email{$^3$joaopsantos@unb.br}

\keywords{Ricci solitons, Harmonic Weyl tensor, Multiply warped products.}
\subjclass[2020]{
53C20, % Global Riem geom
 53E20, % Ricci flows
 53C21, % Methods of Riem. geom. including PDE methods
 53C25, % Einstein manifolds
}
\maketitle{}

\begin{abstract}
In this article, we give a new proof of a result due to J. Kim, which states that the Ricci tensor of a gradient Ricci soliton with dimension $n \geq 4$ and harmonic Weyl tensor has at most three distinct eigenvalues. This result constitutes an essential step in the classification of such manifolds, originally established by J. Kim in dimension $4$ and subsequently extended to dimensions $n\geq5$. Our proof offers two notable advantages: it is shorter and does not require the use of any specialized moving frame.
\end{abstract}
\maketitle

\section{Introduction and main results}
 
\hspace{.5cm}
A gradient Ricci soliton $(M, g, f, \lambda)$ is a four-tuple composed by a Riemannian manifold $(M^{n},g)$, with $n\geq3$, a smooth function $f\in C^{\infty}(M)$ and a constant $\lambda\in\mathbb{R}$ satisfying
\begin{align}\label{fundeq}
	\mathrm{Ric} +  \nabla^2 f= \lambda g.
\end{align}

These metrics have gained prominence due to their deep connection with the Ricci flow. Gradient Ricci solitons not only provide special self-similar solutions to the flow, but also frequently arise as singularity models, appearing as limits of dilations near singularities. From a dynamical perspective, they may be regarded as fixed points of the flow, modulo rescaling. Consequently, they play a central role in revealing the local and asymptotic geometry of solutions to the Ricci flow. As Catino and Mastrolia remark \cite{catinoBook}, substantial progress has been made in classifying such manifolds in dimension three, particularly with the shrinking case being completely resolved and with significant advances for the steady and expanding cases. Despite these results, a comprehensive understanding, even in three dimensions, remains elusive. This challenge becomes even more pronounced in higher dimensions, providing strong motivation for continued investigation. It is then natural to seek classifications of gradient Ricci solitons under additional curvature conditions that may enforce rigidity.

In this direction, the Einstein condition serves as a compelling first step: taking $f$ constant produces the trivial solitons $\mathrm{Ric}=\lambda g$. It is well known that Einstein manifolds have constant sectional curvature if either $n=3$, or $n\geq4$ and the Weyl tensor vanishes; the latter being equivalent to local conformal flatness. Complete gradient Ricci solitons have been classified under the assumptions that either $n=3$ and $\lambda>0$, or $n\geq4$, local conformal flatness, and $\lambda\geq0$ (see \cite{cachen,caoqian,cccmm} and references therein). Moreover, every Einstein manifold has harmonic Weyl curvature, since the Cotton tensor vanishes in the Einstein case. In this context, manifolds with harmonic Weyl curvature arise as natural generalizations of both the Einstein condition and local conformal flatness. The classification of complete gradient Ricci solitons with harmonic Weyl tensor and $\lambda>0$ was obtained by combining the results of \cite{fega} and \cite{muse}. When $\lambda=0$ and $n=4$, these solitons were classified in \cite{kim1}. More recently, the same author achieved a classification in \cite{kim3} for $\lambda\geq0$ and $n\geq5$, recovering the results of \cite{fega} and \cite{muse} for $\lambda > 0$, and also providing a local representation as a multiply warped product for any constant $\lambda\in\mathbb{R}$ and $n\geq4$. These manifolds were further studied in \cite{feng}, where a classification was obtained under additional hypotheses.

A central aspect of Kim’s work \cite{kim3} is the bound on the number of distinct eigenvalues of the Ricci tensor of a gradient Ricci soliton with harmonic Weyl curvature: there can be at most three distinct eigenvalues at each point; more precisely, using a basis of eigenvectors \(\{E_i\}_{i\ge1}\), with $E_1=\nabla f/\vert\nabla f\vert$, the eigenvalues \(\{\lambda_i\}_{i\ge2}\) take at most two distinct values. This was proved for \(n=4\) in \cite{kim1} and extended to \(n\ge5\) in \cite{kim3}. This bound is crucial for several reasons:
\begin{itemize}
	\item The soliton equation and the harmonicity of the Weyl tensor, which are tensorial PDEs, reduce to a system of three ODEs for the potential function and at most two warping data.
 \item It forces the tangent bundle of each level hypersurface of \(f\) to split into at most two totally umbilical distributions, ensuring that \(M\) is locally a multiply warped product with at most two fibers.
 \item By analyzing the local models and using the real analyticity of \(g\) and \(f\) to guarantee smooth transitions between regions, one obtains the global classification by gluing together the local pieces. This approach would become substantially more complicated if four or more eigenvalues were possible.
    
\end{itemize}

Kim took these steps in \cite{kim1} to obtain the classification when $n=4$. As remarked in \cite{kim3}, he used an exhaustive case-by-case analysis of connection components in dimension $4$. It is also mentioned that a reasoning similar to the one used in \cite{kim1} becomes impractical as the dimension increases, because the number of connection coefficients grows significantly. Thus, he introduces in \cite{kim3} a refined frame field $\{F_i\}_{i=1}^n$, where $F_{1}=E_{1}$. This frame is obtained via the parallel transport of an initial ordinary orthonormal eigenframe along the $E_1$ flow, which satisfies $\nabla_{F_1}F_\ell=0$ for $\ell>1$. After a few lengthy computations, this refinement provides just enough control on the Riemannian connection to push his approach through in higher dimensions. Subsequently, assuming at least three distinct $\lambda_i$ for $i\ge 2$, Kim obtains contradictory identities, using an $s$-invariant quantity defined using ODEs which arise naturally in the analysis; thus, more than three Ricci eigenvalues cannot occur. This ensures the metric can be written as a multiply warped product with (at most) two Einstein fibers.

The main goal of this article is to provide a new proof of the estimate on the number of distinct eigenvalues, mentioned above. Namely, we give an alternative and shorter proof of the following theorem:

\begin{theorem}[Kim]\label{maxxnumeig-INTRO}
Let $(M^n,g,f,\lambda)$, $n\geq4$, be a gradient Ricci soliton with harmonic Weyl curvature and nonconstant $f$. Then, the Ricci tensor of $M$ has at most three distinct eigenvalues at each point of $M$.
\end{theorem}

We now outline our proof of Theorem \ref{maxxnumeig-INTRO}, which deviates from Kim’s methods in a few significant aspects. Our approach begins by establishing the local geometric structure of the soliton {\it before} addressing the bound on the number of distinct eigenvalues of the Ricci tensor. Thus, we first obtain its local decomposition as a multiply warped product
\begin{align*}
I \times_{h_1} N_1^{r_1} \times \dots \times_{h_k} N_k^{r_k},    
\end{align*}
with warping functions $h_i$, for $i\in\{1,\ldots,k\}$, in accordance with the requirements of Remark \ref{grmnt_wrpngfnctn}. Then we construct a nonconstant polynomial of degree at most two that has $\xi_{i}=h'_i/h_i$ as roots, showing that $k\leq2$. Combining the last fact with a relation between the $\xi_i$ and the eigenvalues of the Ricci tensor, we obtain the desired result.

A key observation is that neither the local geometric decomposition nor the estimate on $k$ is achieved by resorting to any specialized moving frames. Instead, we exploit the integrability of the distributions generated by the eigenspaces of $\mathrm{Ric}$ to introduce a natural system of local coordinates. Namely, we use the arc-length parameter $s$ along the integral curves of $\nabla f/\vert\nabla f \vert=E_{1}$ and choose coordinates on each integral manifold of the distributions associated to the Ricci tensor. Using these coordinates, we prove the multiply warped product structure (see Lemma \ref{lemmamany}) and construct the polynomial (see Lemma \ref{lem_ply_deg2_}).

\section{Preliminaries}\label{prelim}

\hspace{.5cm}Consider vector fields $X,Y,Z,T\in\mathfrak{X}(M)$.  We will adopt the following convention for the curvature
\begin{align*}
	\mathrm{Rm}(X,Y,Z,T)=\left\langle\nabla_{Y}\nabla_{X}Z-\nabla_{X}\nabla_{Y}Z+\nabla_{[X,Y]}Z,T\right\rangle.
\end{align*}
For the Weyl and Cotton tensors, recall that they are defined, respectively, as
\begin{align}
	W(X,Y,Z,T)=&\mathrm{Rm}(X,Y,Z,T)-\frac{1}{n-2}\left(\left(\mathrm{Ric}-\frac{R}{n}g\right)\varowedge g\right)(X,Y,Z,T)\nonumber\\
    &-\frac{R}{2n(n-1)}(g\varowedge g)(X,Y,Z,T) \nonumber,\\
	C(X, Y, Z) =&\left(\nabla_X \mathrm{Ric}\right)(Y, Z) -  \left(\nabla_Y \mathrm{Ric}\right)(X, Z)\label{cotton}\\
    &-\frac{1}{2(n-1)} \left\{\left(\nabla_X (Rg)\right)(Y, Z) - \left(\nabla_Y (Rg)\right)(X, Z)\right\}.\nonumber
\end{align}
where $\varowedge $ is the Kulkarni-Nomizu product, whose definition is given in \cite[page 47]{besse}, for example.

Recall that a Riemannian manifold is locally conformally flat if and only if either its Weyl tensor vanishes and $n\geq4$, or its Cotton tensor vanishes and $n=3$. Furthermore, the Weyl tensor always vanishes in dimension three, and the Weyl tensor is harmonic if and only if the Cotton tensor vanishes and $n\geq4$.

\subsection{Results for when the Cotton tensor vanishes}

\hspace{.5cm}We start by recalling results from \cite{fega} and \cite{kim1}, useful in the rest of this paper. In the version of Lemma 2.1 below, we emphasize the fact that it is an equivalence.

\begin{lemma}[\cite{fega}]\label{bari}
	Suppose $(M^n, g, f, \lambda)$ is a gradient Ricci soliton. Then $M$ has zero Cotton tensor if, and only if,
	\begin{align*}
		\mathrm{Rm}(\nabla f,X,Y,Z)&=Y\left(\frac{R}{2(n-1)}\right)g(X,Z)-Z\left(\frac{R}{2(n-1)}\right)g(X,Y)\nonumber\\
		&=\frac{1}{n-1}(\mathrm{Ric}(\nabla f,Y)g(X,Z)-\mathrm{Ric}(\nabla f,Z)g(X,Y))
	\end{align*}
	for all vector fields $X,Y,Z \in \mathfrak{X}(M)$. If, in particular, $n\geq4$, the condition above is equivalent to the soliton having harmonic Weyl curvature.
\end{lemma}

\begin{proof}
	It is easy to see that on a gradient Ricci soliton we have
	\begin{align*}
		\left(\nabla_X \mathrm{Ric}\right)(Y, Z)
		&=g \left( \nabla_{\nabla_{X} Y } \nabla f - \nabla_{X} \nabla_Y \nabla f, Z    \right)
	\end{align*}
	Consequently,
	\begin{align*}
		\left( \nabla_{X} \mathrm{Ric}  \right) \left( Y, Z \right)-\left( \nabla_{Y} \mathrm{Ric}  \right) \left( X, Z \right) = \mathrm{Rm} \left( X, Y,\nabla f,Z \right).
	\end{align*} 
	Straightforward computations also yield
	\begin{align}
			 \left( \nabla_{X} \left(Rg \right)   \right) \left( Y,Z \right) - \left( \nabla_{Y} \left(Rg\right)   \right) \left( X,Z \right) =X(R) g \left( Y,Z \right) - Y(R) g(X,Z).
	\end{align}
	Therefore, using \eqref{cotton}, the Cotton tensor $C$ vanishes if and only if
	\begin{align*}
			\mathrm{Rm} \left( \nabla f, Z,Y,X \right) = Y \left( \frac{R}{2 \left( n-1 \right) } \right) g(X, Z) - X \left( \frac{R}{2(n-1)}  \right) g \left( Y,Z \right)  .
	\end{align*}
	This last expression is equivalent to the first equation we aimed to prove. The second equality follows directly from identities in \cite[page 462]{fega}.
\end{proof}

\begin{lemma}[Lemma 2.2 of \cite{fega}]\label{caochen}
	Let $(M,g,f,\lambda)$ be a gradient Ricci soliton with zero Cotton tensor and nonconstant $f$. Let $c$ be a regular value of $f$ and $\Sigma_{c}=f^{-1}(c)$ be the level surface of $f$. Then,
	\begin{enumerate}
		\item Where $\nabla f\neq0$, $E_{1}=\frac{\nabla f}{|\nabla f|}$ is an eigenvector of $\mathrm{Ric}$.
		\item\label{cao2} $|\nabla f|$ is constant on a connected component of $\Sigma_{c}$.
		\item\label{cao3} There is a function $s$ locally defined with $s(x)=\int\frac{\ \mathrm{d} f}{|\nabla f|}$, so that $\ \mathrm{d} s=\frac{\ \mathrm{d} f}{|\nabla f|}$ and $E_{1}=\nabla s$.
		\item\label{cao4} $E_{1}E_{1}f=-\mathrm{Ric}(E_{1},E_{1})+\lambda$. In particular, $\lambda_1 = \mathrm{Ric}(E_{1},E_{1})$ is constant on a connected component of $\Sigma_{c}$.
		\item\label{cao5} Near a point in $\Sigma_{c}$, the metric $g$ can be written as
		\begin{align*}
			g=\ \mathrm{d} s^2+\sum_{i,j\geq2}g_{ij}(s,x_{2},\ldots,x_{n}) \ \mathrm{d} x_{i}\otimes  \mathrm{d} x_{j}.
		\end{align*}
		\item $\nabla_{E_{1}}E_{1}=0$.
	\end{enumerate}
\end{lemma}

It is a well-known fact that a Riemannian manifold $(M^n, g)$, $n\geq4$, has a harmonic Weyl tensor if and only if its Schouten tensor $\mathcal{A}=\mathrm{Ric}-\frac{R}{2(n-1)}g$ is Codazzi. In coordinates, this is equivalent to
\begin{align}
	\nabla_{i}\mathcal{A}_{jk}=\nabla_{j}\mathcal{A}_{ik}.
\end{align}
Let $\mathcal{A}$ be a Codazzi tensor and denote by $E_{\mathcal{A}}(x)$ the number of distinct eigenvalues of $\mathcal{A}$ at $x$. In \cite{derd}, Derdzinski considered the following open dense set
\begin{align}
	M_{\mathcal{A}}=\{x\in M \ |  \ E_{\mathcal{A}}(x)\text{ is constant in a neighborhood of }x\}.
\end{align}
It turns out that in $M_{\mathcal{A}}$ the eigenvalues of $\mathcal{A}$ are well-defined and define smooth functions $\lambda_{1},\ldots,\lambda_{n}:M_{\mathcal{A}}\rightarrow\mathbb{R}$. Furthermore, he proved that in such a set the following is true\\

\begin{lemma}[Derdzi\'{n}ski]\label{derdlemma}
	Let $(M^n,g)$, $n\geq4$, be a Riemannian metric with harmonic Weyl curvature. Let $\{E_{i}\}_{i=1}^{n}$ be a local orthonormal frame such that $\mathrm{Ric}(E_{i},\cdot)=\lambda_{i}g(E_{i},\cdot)$. Then,
	\begin{enumerate}
		\item\label{derd1} For any $i,j,k\geq1$,
		\begin{align*}
			&(\lambda_{j}-\lambda_{k})\left\langle\nabla_{E_{i}}E_{j},E_{k}\right\rangle+\nabla_{E_{i}}(\mathcal{A}(E_{j},E_{k}))=\\
			&(\lambda_{i}-\lambda_{k})\left\langle\nabla_{E_{j}}E_{i},E_{k}\right\rangle+\nabla_{E_{j}}(\mathcal{A}(E_{k},E_{i})).
		\end{align*}
		\item If $k\neq i$ and $k\neq j$, then $(\lambda_{j}-\lambda_{k})\left\langle\nabla_{E_{i}}E_{j},E_{k}\right\rangle=(\lambda_{i}-\lambda_{k})\left\langle\nabla_{E_{j}}E_{i},E_{k}\right\rangle$.
		\item Given distinct eigenfunctions $\lambda$ and $\mu$ of $\mathcal{A}$ and local vector fields $U$ and $V$ such that $A(V)=\lambda V$ and $A(U)=\mu U$ with $|U|=1$, it holds that $V(\mu)=(\mu-\lambda)\left\langle\nabla_{U}U,V\right\rangle$.
		\item Each distribution $D_{\lambda_{i}}$, defined by $D_{\lambda_{i}}(p)=\{v\in T_{p} M \ | \ \mathrm{Ric}(v,\cdot)=\lambda_{i}g(v,\cdot)\}$, is integrable and its leaves are totally umbilical submanifolds of $M$.
	\end{enumerate}
\end{lemma}

\subsection{Multiply warped products}

\hspace{.5cm}In this section, we collect some formulas for the curvatures of a multiply warped product
\begin{align}\label{MWPS}
	M = B \times_{h_1} N_1^{r_1} \times \cdots \times_{h_k} N_{k}^{r_k}.
\end{align}
Recall that \eqref{MWPS} means that the manifold $B\times N_1^{r_1} \times \cdots \times N_{k}^{r_k}$ is endowed with the Riemannian metric
\begin{align}\label{metricMWPS}
	g=g_{B}+h_1^2 g_{N_1}+\cdots+h_{k}^2 g_{N_k}.
\end{align}
In this context, $B$ is called the base and $N_{1},\ldots,N_{k}$ the fibers of the multiply warped product; each positive smooth function $h_{i}:B\rightarrow\mathbb{R}$, $i\in\{1,\cdots,k\}$, is called the warping function corresponding to the fiber $N_{i}^{r_{i}}$. When $k=1$, this is simply called warped product, and a classical reference on it is \cite{oneill}.

	\begin{remark}[Number of fibers]\label{grmnt_wrpngfnctn}
	We adopt the following conventions, which are similar to those of \cite{brovaz}. Namely,
\begin{enumerate}[label=(\roman*)]
\item any fiber whose warping function is constant is absorbed into the base; and
\item if two warping functions differ by a positive constant factor, we rescale the corresponding fiber metrics and identify the fibers.
\end{enumerate}
Consequently, all warping functions are nonconstant and pairwise nonproportional. The \textbf{number of fibers} is the integer $k$ given by the number of equivalence classes of nonconstant warping functions under $f \sim c\,g$ for $c>0$.
\end{remark}

Metrics such as \eqref{metricMWPS} have been used to give examples of manifolds with additional geometric properties \cite{bishop,brovaz,choui,danwan} and arise naturally in a variety of circumstances \cite{caoqian,cccmm,caozhou,kim1,kim3,feng,shin}.

There are a couple of works describing the geometry of these metrics, and here we refer to \cite{bishop,brovaz,DobUn,oneill}. The next lemma collects the Levi-Civita connection and the Ricci tensor of a multiply warped product in terms of the corresponding ones on the base, the fibers, and quantities associated with the warping functions \cite[Proposition 2.5, Proposition 2.6]{DobUn}.

\begin{lemma}[\cite{DobUn}]\label{warped}
	Let $M = B\times_{h_1} N_1^{r_1} \times \cdots \times_{h_k} N_{k}^{r_k}$ be a multiply warped product, and consider $X, Y, Z \in \mathcal{L}(B)$, $V \in \mathcal{L}(N_i)$ and $W \in \mathcal{L}(N_j)$ lifted vector fields. Then:
	
	\begin{enumerate}
		\item The covariant derivative of $M$ satisfies the following relations
		\begin{align*}
			&\nabla_X Y =\nabla^{B}_X Y ,\\
			&\nabla_X V = \nabla_V X = \frac{X(h_i)}{h_i} V,\\
			&\nabla_W V = 0,\ \ \text{if}\ \ i\neq j,\\
			&\nabla_W V = \nabla^{N_{i}}_W V-h_{i}g_{N_{i}}(W,V)\nabla^{B}h_{i},\ \ \text{if}\ \ i=j.
		\end{align*}
		
		\item The Ricci tensor of $M$ is given by
		\begin{align*}
			&\mathrm{Ric}(X, Y) = \mathrm{Ric}_{B}(X,Y) - \sum_{1 \leq \ell \leq k} \frac{r_{\ell}}{h_{\ell}} \nabla^{2}_{B}h_{\ell}(X, Y),\\
			&\mathrm{Ric}(X, V) = 0,\\
			&\mathrm{Ric}(V, W) = 0, \text{ if } i \neq j,\\
			&\mathrm{Ric}(V, W) = \mathrm{Ric}_{N_i}(V, W) - \left(\frac{\Delta_B h_i}{h_i} + (r_i - 1) \frac{| \nabla_{B} h_i |^2}{h_i^2} \right.\label{RicVWi}\\ &\quad\quad\quad\quad\quad + \left.\sum_{\substack{1 \leq \ell \leq k \\
					\ell \neq i  \\
			}} r_{\ell} \frac{g_{B}(\nabla_{B} h_i, \nabla_{B} h_{\ell})}{h_i h_{\ell}} \right) g(V, W), \text{ if } i = j.\nonumber
		\end{align*}
		
		 \item The scalar curvature of $M$ is given by
		 \begin{equation*}
		 	\begin{aligned}
		 		R = &R_{B} - 2 \sum_{1 \leq i \leq k} r_i \frac{\Delta_{B} h_i}{h_i} + \sum_{1 \leq i \leq k} \frac{R_{N_i}}{h_i^2} - \sum_{1 \leq i \leq k} r_i(r_i - 1) \frac{| \nabla_{B} h_i |^2}{h_i^2}\\&-\sum_{1 \leq i \leq k} \sum_{\substack{1 \leq \ell \leq k \\
		 				\ell \neq i  \\
		 		}} r_i r_{\ell} \frac{g_{B}(\nabla_{B} h_i, \nabla_{B} h_{\ell})}{h_i h_{\ell}}.
		 	\end{aligned}
		 \end{equation*}
	\end{enumerate}
\end{lemma}

\section{Local representation as a multiple warped product} \label{cotton_results}
\hspace{.5cm}Let $(M,g,f,\lambda)$ be a gradient Ricci soliton with harmonic Weyl tensor. The goal of this section is to give a new proof of the following result, established by Kim in \cite{kim3}
\begin{theorem}\label{decompwarp_lmbd_ctt}
Any gradient Ricci soliton with harmonic Weyl curvature is locally a multiply warped product $I\times_{h_1} N_1^{r_1}\times\cdots\times_{h_k} N_{k}^{r_k}$ of a one-dimensional base $I$ and $k$ fibers $N_{i}^{r_{i}}$. Furthermore, each fiber $N_{i}^{r_{i}}$ of dimension $r_{i}\geq2$ must be Einstein.
\end{theorem}

The first step, presented in the next subsection, shows that certain functions depend only on the arc length of the integral curve of $\frac{\nabla f}{\vert\nabla f\vert}$. In the subsequent subsection, we compute certain components of the metric using an appropriate coordinate system, which allows us to conclude the local decomposition as a multiply warped product. This avoids the necessity of using a special moving frame.

\subsection{$R$ and the Eigenvalues of $\mathrm{Ric}$ depend only on $s$}\label{subsec_dep_s}

\hspace{.5cm}In order to prove the local representation of $M$ as a multiple warped product, we first show that using the system of coordinates of item \eqref{cao5} of Lemma \ref{caochen}, certain important functions depend only on $s$, the arc length of the integral curve of $\frac{\nabla f}{\vert\nabla f\vert}$.

First, we deal with the set $\mathcal{R}$ of regular points of the potential function $f$ of the gradient Ricci soliton $M$, that is,
\begin{align*}
	\mathcal{R}=\{x\in M \ | \ \nabla f(x)\neq0\}.
\end{align*}
Since a gradient Ricci soliton is real analytic in harmonic coordinates (\cite{ivey2}), $\mathcal{R}$ is dense in $M$.

For each point $p$ of the open and dense subset $\mathcal{R}\cap M_{\mathcal{A}}$, we will consider the orthonormal frame $\{E_{i}\}_{i=1}^n$ given in Lemma \ref{derdlemma} and recall that $E_{1}=\frac{\nabla f}{|\nabla f|}$. For this frame, we have $\mathrm{Ric}_{ij}=\lambda_{i}\delta_{ij}$. Furthermore,
\begin{lemma}\label{nablaframe}
	Let $(M^n,g,f,\lambda)$, $n\geq4$, be a gradient Ricci soliton with harmonic Weyl curvature and nonconstant $f$. Then,
	\begin{align}\label{step2}
		\nabla_{E_{a}}E_{1}=\xi_{a}E_{a}\ \ \text{and}\ \ \xi_{a}=-\left\langle\nabla_{E_{a}}E_{a},E_{1}\right\rangle, \ \forall a \geq 2,
	\end{align}
	where
	\begin{align}\label{step1}
		\xi_{a}=\frac{\lambda-\lambda_{a}}{|\nabla f|}.
	\end{align}
\end{lemma}
\begin{proof}
	To prove the first identity of $(\ref{step2})$, notice that for any $a\geq2$ we have
	\begin{align}\label{comp1}
		\begin{split}
			\nabla_{E_{a}}E_{1}=\frac{\nabla_{E_{a}}\nabla f}{|\nabla f|}=\frac{\lambda E_{a}-\mathrm{Ric}(E_{a},\cdot)}{|\nabla f|}=\frac{(\lambda-\lambda_{a})E_{a}}{|\nabla f|}=\xi_{a}E_{a}.
		\end{split}
	\end{align}
	Now we combine it with the equality $\left\langle\nabla_{E_{a}}E_{1},E_{a}\right\rangle=-\left\langle\nabla_{E_{a}}E_{a},E_{1}\right\rangle$ to get the second identity of $(\ref{step2})$.
	
\end{proof}

It follows from Lemma \ref{caochen} that the scalar curvature $R$ and the first Ricci-eigenvalue $\lambda_1$ are constant on each connected component of a regular level set $\Sigma_{c}$ of $f$. The following lemma ensures the same holds for all the remaining eigenvalues $\lambda_i, i \geq 2$. Consequently, these quantities depend solely on the arc-length parameter $s$ (cf. item \eqref{cao5} of Lemma \ref{caochen}). This result was originally established in \cite{kim1} for four-dimensional Ricci solitons, and as Kim remarks in \cite{kim3}, one can easily see that the arguments can be naturally extended to higher dimensions.

\begin{lemma}
	Let $(M^n,g,f,\lambda)$, $n\geq4$, be a gradient Ricci soliton with harmonic Weyl curvature and nonconstant $f$. The functions $\lambda_{2},\ldots,\lambda_{n}$ are constant on each connected component of $\Sigma_{c}$. As a consequence, the functions $\xi_{2},\ldots,\xi_{n}$ are also constant on each connected component of $\Sigma_{c}$.
\end{lemma}

\subsection{Local multiply warped product structure}\label{der_ivey_ex}

\hspace{.5cm}In this subsection, we prove that gradient Ricci solitons with zero Cotton tensor are multiply warped products around regular points of the potential function.

Recall that Derdziński's result (cf. Lemma \ref{derdlemma}) ensures the distributions $D_{\lambda_i} = \mathrm{Span}\{E_{\ell} \ \vert \ \ell \in [i] \}$ are integrable. Denote by $M_{i}^{r_i}$ the integral manifolds of $D_{\lambda_i}$. We also adopt the following notation, following \cite{feng}: for each $a\geq2$ we consider the set of indices
\begin{align}
	[a]=\{j\in\{2,\ldots,n\} \ | \ \lambda_{j}=\lambda_{a}\}.
\end{align}
We also adopt the convention that $2\leq a,b,c,\ldots,\alpha,\beta,\gamma,\ldots\leq n$ satisfy $b,c\in[a]$, $\beta,\gamma\in[\alpha]$ and $[a]\neq[\alpha]$.

Once we know the distributions corresponding to the eigenspaces of the Ricci tensor are integrable, we can make use of coordinate systems. This will make some computations simpler than those performed in \cite{kim1,feng,shin}. \\

\begin{lemma}\label{lemmamany}
	Let $(M^n,g,f,\lambda)$, $n\geq4$, be a gradient Ricci soliton with harmonic Weyl curvature and nonconstant $f$. Let $(x_{b})_{b\in[a]}$ and $(x_{\beta})_{\beta\in[\alpha]}$ be local coordinate systems of the integral manifolds $M^{r_a}_{a}$ and $M^{r_\alpha}_{\alpha}$ of the distributions $D_{a}$ and $D_{\alpha}$, respectively. Setting $\partial_{1}=E_{1}$, we have
	\begin{align}\label{37lemma36}
		\partial_{1}g_{ab}=2\xi_{a}g_{ab}\ \ \ \ \text{and}\ \ \ \ \ \partial_\alpha g_{ab}=0.
	\end{align}
\end{lemma}
\begin{proof}
    In order to prove the lemma, we will first show that
	\begin{align}\label{three_id}
	    \nabla_{\partial_{a}}{\partial_{1}}=\xi_{a}\partial_{a},\ \ \text{and}\ \ \ \nabla_{\partial_{\alpha}}\partial_{a}=0.
	\end{align}
        The first equality of \eqref{three_id} follows from considering any vector field $X$ and $a\geq2$, and noticing that
	\begin{align*}
		\left\langle\nabla_{\partial_a}\partial_{1},X\right\rangle=\frac{1}{|\nabla f|}\nabla^2 f(\partial_a,X)=\frac{\lambda-\lambda_{a}}{|\nabla f|}\left\langle\partial_{a},X\right\rangle=\left\langle\xi_{a}\partial_{a},X\right\rangle.
	\end{align*}
        To prove the second equality of \eqref{three_id}, let $a,\alpha,z\in\{2,\ldots,n\}$ are so that $[a]\neq[\alpha]$, then
	\begin{align*}
		\mathrm{Rm}_{a\alpha1z}=&\left\langle\nabla_{\partial_{\alpha}}\nabla_{\partial_{a}}\partial_{1}-\nabla_{\partial_{a}}\nabla_{\partial_{\alpha}}\partial_{1},\partial_{z}\right\rangle\\
		=&\xi_{a}\left\langle\nabla_{\partial_{\alpha}}\partial_{a},\partial_{z}\right\rangle-\xi_{\alpha}\left\langle\nabla_{\partial_{a}}\partial_{\alpha},\partial_{z}\right\rangle\\
		=&(\xi_{a}-\xi_{\alpha})\left\langle\nabla_{\partial_{\alpha}}\partial_{a},\partial_{z}\right\rangle.
	\end{align*}
	On the other hand, by using $(\ref{bari})$ we obtain
	\begin{align*}
		(\xi_{a}-\xi_{\alpha})\left\langle\nabla_{\partial_{\alpha}}\partial_{a},\partial_{z}\right\rangle=\mathrm{Rm}_{1za\alpha}
		=\partial_{a}\left(\frac{R}{2(n-1)}\right)g_{z\alpha}-\partial_{\alpha}\left(\frac{R}{2(n-1)}\right)g_{za}
		=0,
	\end{align*}
	where in the last equality we have used that $\partial_{j}R=0$, for all $j\geq2$. As $[a]\neq[\alpha]$ and $z\in\{1,\ldots,n\}$, we conclude that $\nabla_{\partial_{\alpha}}\partial_{a}=0$.

	In order to finish the proof of the lemma, recall that $\xi_{a}=\xi_{b}$ and consider the following derivatives of $g_{ab}$
	\begin{align*}
		\partial_{1}g_{ab}&=g(\nabla_{\partial_{1}}\partial_{a},\partial_{b})+g(\partial_{a},\nabla_{\partial_{1}}\partial_{b})=2\xi_{a}g_{ab}\\
		\partial_{\alpha}g_{ab}&=g(\nabla_{\partial_{\alpha}}\partial_{a},\partial_{b})+g(\partial_{a},\nabla_{\partial_{\alpha}}\partial_{b})=0.
	\end{align*}
\end{proof}

In what follows, we use Lemma \ref{lemmamany} to prove the local decomposition of the metric as a multiply warped product.

\begin{theorem}\label{decompwarp}
Any gradient Ricci soliton with harmonic Weyl curvature is locally a multiply warped product $I\times_{h_1} N_1^{r_1}\times\cdots\times_{h_k} N_{k}^{r_k}$ of a one-dimensional base $I$ and $k$ fibers $N_{i}^{r_{i}}$. Furthermore, each fiber $N_{i}^{r_{i}}$ of dimension $r_{i}\geq2$ must be Einstein.
\end{theorem}

\begin{proof}
	Let $U \subset M_{\mathcal{A}} \cap \{ \nabla f \neq 0 \}$ be an open connected subset, and consider a system of local coordinates $(x_1 = s, x_2, x_3, \ldots, x_n)$ in $U$ as in Lemma \ref{lemmamany}, so that $U=I\times N_1^{r_1}\times\cdots\times N_{k}^{r_k}$, topologically. Let us fix $i\in\{1,\ldots,k\}$, $s_0 \in I$, $(x_{1}^{0}=s_0,x_{2}^{0},\ldots,x^{0}_{n})\in U$ and $a, b \in [i]$. \emph{Mutatis mutandis}, equation \eqref{37lemma36} from Lemma \ref{lemmamany} implies that 
	\[
	\partial_1g_{ab} = 2 \xi_i g_{ab},
	\]
	where we are using that $\xi_{i}=\xi_{a}$. Therefore, if we define the function $\tilde{h}_i: I \to \mathbb{R}$ by
	\[
	s \in I \mapsto \tilde{h}_i(s) := \mathrm{exp}\left(\int_{s_0}^s \xi_i(y) \ \mathrm{d}y \right)
	\]
	Then $\tilde{h}_i$ satisfies
	\begin{align*}
	    \partial_1\left(\tilde{h}_i^{-2} g_{ab} \right) = 0\ \ \ \ \text{and}\ \ \ \ \partial_\alpha\left(\tilde{h}_i^{-2} g_{ab} \right) = 0,
	\end{align*}
	so that $\tilde{h}^{-2}_i g_{ab}$ depends only on $(x_{b})_{b\in[a]}$. In particular, it is constant in $s$, i.e, 
	\[
	(\tilde{h}_i(s))^{-2} g_{ab}(s, x_2, \ldots, x_n) = (\tilde{h}_i(s_0))^{-2} g_{ab}(s_0, x_2, \ldots, x_n)
	\]
	or, equivalently,
	\begin{equation*}
		g_{ab}(s, x_2, \ldots, x_n) = h_{i}(s)^2g_{ab}(s_0, x_2, \ldots, x_n)
	\end{equation*}
	where $h_{i}(s) := \frac{\tilde{h}_i(s)}{\tilde{h}_i(s_0)}$. Without any loss of generality, we can assume $h_i(s_0)=1$. This gives $g_{ab}= h_{i}^2(g_{N_{i}})_{ab}$, where $g_{N_{i}}$ is a metric in $N_{i}^{r_{i}}$. In an entirely analogous manner, we obtain Riemannian metrics $g_{\alpha}$ on $N_{\alpha}$ for any $\alpha \notin [i]$. This proves that in $U$, $g$ can be written as the following (multiply) warped product
	\begin{align*}
		g = \ \mathrm{d}s^2 + h_i^2 g_i + \sum_{\alpha \notin [i]} h_{\alpha}^2 g_{\alpha}.
	\end{align*}
	
	Now we prove that the fibers $N_{i}^{r_{i}}$ of dimension $r_{i}\geq2$ must be Einstein. To see this, consider $V, W \in \mathcal{L}(N_{i})$ and notice that combining \eqref{fundeq} and the equations for the covariant derivative and the Ricci tensor, given in Lemma \ref{warped}, we have the following equality
	\begin{align*}
		\mathrm{Ric}_{N_{i}}(V,W)= \left(h_ih''_i + (r_i - 1)(h'_i)^2+ h_ih'_i \displaystyle \sum_{\alpha \notin [i]} r_{\alpha} \frac{h'_{\alpha}}{h_{\alpha}}+h_ih'_if'+\lambda h_i^2 \right) g_{N_{i}}(V, W).
	\end{align*}
	Now notice that the function multiplying $g_{N_{i}}$ depends only on $s$, while both $\mathrm{Ric}_{N_{i}}$ and $g_{N_{i}}$ depend only on $(x_2, \ldots, x_n)$. This shows that there is a constant $\mu_{i}$ so that $\mathrm{Ric}_{N_{i}}=\mu_{i}g_{N_{i}}$, as claimed. A similar argument shows that each $N_{\alpha}$ with $r_{\alpha}\geq2$ is Einstein.
\end{proof}

A straightforward consequence of the proof of Theorem \ref{decompwarp} gives the following important identity.
\begin{corollary} Assuming the hypothesis of Theorem \ref{decompwarp}, and the notation of Lemma \ref{nablaframe}, we have 
	\begin{equation}\label{eqXiRelFuncWarp}
		\frac{\lambda - \lambda_i}{f'} = \xi_i = \frac{h_{i}'}{h_i}.
	\end{equation}
\end{corollary}

\section{Estimating the number of fibers}\label{number_fiber}

\hspace{.5cm}In this section, we prove that the number of fibers in the representation given in Theorem \ref{decompwarp} is at most two. More precisely, we prove that:
\begin{theorem}\label{decompwarpint}
Any nontrivial gradient Ricci soliton with harmonic Weyl curvature is locally a multiply warped product with a one-dimensional base and at most two fibers. Furthermore, the fibers of dimension at least two must be Einstein.
\end{theorem}

As we will see below, the theorem above already implies that the Ricci tensor of $M$ has at most three distinct eigenvalues, which proves Theorem \ref{maxxnumeig-INTRO}. 

 To prove Theorem \ref{decompwarpint}, we will show that the quantities $h_{i}'/h_{i}$ satisfy a nonconstant polynomial of degree at most two. In order to construct this polynomial, we will use the following lemma proved by F. Li in \cite{feng}.

\begin{lemma}[Li, F.]\label{lema38}
	Let $(M^n,g,f,\lambda)$, $n\geq4$, be a gradient Ricci soliton with harmonic Weyl curvature and nonconstant $f$. Then the following equations hold
	\begin{align}
		&\xi'_{i}+\xi^{2}_{i}=-\frac{R'}{2(n-1)f'}\label{excep}\\
		&-f^{\prime} \xi_i+\lambda=-\xi_i^{\prime}-\xi_i \sum_{j=2}^n \xi_j+(r_i-1) \frac{\mu_i}{h_i^2}\label{sec38}\\
		&\lambda_1=-f^{\prime \prime}+\lambda=-(n-1)\left(\xi_i^{\prime}+\xi_i^2\right)\label{third38}
	\end{align}
\end{lemma}

It will be convenient to introduce the functions
    \begin{align}\label{def_B_C}
        B=\frac{(n-1)\lambda-R+\lambda_{1}-(f')^2}{f'}\ \ \ \text{and}\ \ \ C=-\frac{R'}{2(n-1)f'}+\lambda.
    \end{align}
With this notation, \eqref{excep} can be written as $\xi_{i}'=-\xi_{i}^{2}+C-\lambda$. Combining this equation, the first equality of \eqref{eqXiRelFuncWarp} and \eqref{sec38}, we obtain

    \begin{align}\label{first_scnd_Dg}
        \xi^{2}_{i}-B\xi_i-C=-(r_i-1) \frac{\mu_i}{h_i^2}.
    \end{align}
This equation allows us to give an alternative proof of \cite[Proposition 3.4]{kim1}, which establishes an estimate to the number of distinct eigenvalues of the Ricci tensor of a gradient Ricci soliton $M^4$ with harmonic Weyl tensor.

\begin{proposition}\label{prop_n=4}
    If $n=4$, then $M$ has at most two fibers in the representation given by Theorem \ref{decompwarp}. Equivalently, the Ricci tensor of $M^4$ has at most three distinct eigenvalues.
\end{proposition}
\begin{proof}
    Assume by contradiction that there are exactly three distinct fibers in the representation given by Theorem \ref{decompwarp}. Then, $\xi_{1},\ \xi_{2}$ and $\xi_{3}$ are pairwise distinct. Furthermore, each fiber must have dimension $1$, that is, $r_{1}=r_{2}=r_{3}=1$. This last information means that \eqref{first_scnd_Dg} simply becomes $\xi^{2}_{i}-B\xi_i-C=0$, for $i\in\{1,2,3\}$. But this implies the existence of at most two distinct $\xi_{i}$, which is a contradiction.
\end{proof}

In what follows, we extend this argument to higher dimensions. Namely, we construct a nonzero polynomial of degree at most two, which has as roots the functions $\xi_{1},\ldots,\xi_{k}$.

\begin{lemma}\label{lem_ply_deg2_}
    The functions $\xi_{1},\ldots,\xi_{k}$ satisfy the following equation
    \begin{align}\label{poly_atmst2}
        B\xi_{i}^2+(B'+2\lambda)\xi_{i}+(C-\lambda)B+C'=0.
    \end{align}
    Furthermore, the polynomial in $\xi_{i}$ defined by the left-hand side of \eqref{poly_atmst2} is nontrivial.
\end{lemma}
\begin{proof}
    Our first step is to eliminate $h_{i}$ from \eqref{first_scnd_Dg}. In order to do this, consider the derivative of this equation, and use the second equality in \eqref{eqXiRelFuncWarp}, to get
    \begin{align*}
        2\xi_{i}\xi'_{i}-B'\xi_i-B\xi_i'-C'=2(r_i-1) \frac{\mu_i}{h_i^2}\xi_{i}=-2(\xi^{2}_{i}-B\xi_i-C)\xi_{i}
    \end{align*}
    Using $\xi_{i}'=-\xi_{i}^{2}+C-\lambda$, we obtain
    \begin{align*}
        2\xi_{i}(-\xi^{2}_{i}+C-\lambda)-B'\xi_i-B(-\xi^{2}_{i}+C-\lambda)-C'=-2(\xi^{2}_{i}-B\xi_i-C)\xi_{i},
    \end{align*}
    which simplifies to \eqref{poly_atmst2}.

    Now, assume that the polynomial defined by the left-hand side of \eqref{poly_atmst2} is trivial. This happens if and only if $B(s)=\lambda=0$, $\forall s\in I$. In what follows, we will show that the vanishing of $B$ and $\lambda$ simultaneously implies that $f$ is constant, which is a contradiction. 
    
    From \eqref{poly_atmst2} and $B=\lambda=0$, we have immediately that $C'=0$. Then, there exists $c_{0}\in\mathbb{R}$ so that
    \begin{align}
        &R'=-2(n-1)c_{0}f',\label{frst_cnsqnc}\\
        &\xi'_{i}=-\xi^{2}_{i}+c_0,\label{excep'}\\
        &f''=-\lambda_1=(n-1)c_{0},\label{third38'}\\
        &\lambda_i=-f'\xi_i\label{eqXiRelFuncWarp''},
    \end{align}
    where these equalities follow from the definition of $C$, \eqref{excep}, \eqref{third38} and \eqref{eqXiRelFuncWarp}, respectively. On the other hand, using $R=\lambda_{1}+\displaystyle\sum_{\ell=1}^{k}r_{\ell}\lambda_{\ell}$, $\vert \mathrm{Ric}\vert^2=\lambda_{1}^{2}+\displaystyle\sum_{\ell=1}^{k}r_{\ell}\lambda_{\ell}^{2}$, \eqref{third38'} and \eqref{eqXiRelFuncWarp''}, we get
    \begin{align}\label{scl_ric}
        R=-(n-1)c_{0}-f'\displaystyle\sum_{\ell=1}^{k}r_{\ell}\xi_{\ell}\ \ \ \ \text{and}\ \ \ \ \vert \mathrm{Ric}\vert^2=(n-1)^{2}c_{0}^{2}+(f')^2\displaystyle\sum_{\ell=1}^{k}r_{\ell}\xi_{\ell}^{2}.
    \end{align}
    Now, this expression for $R$ implies that
    \begin{align*}
        R'&=-f''\displaystyle\sum_{\ell=1}^{k}r_{\ell}\xi_{\ell}-f'\displaystyle\sum_{\ell=1}^{k}r_{\ell}\xi_{\ell}'=-(n-1)c_{0}\displaystyle\sum_{\ell=1}^{k}r_{\ell}\xi_{\ell}+f'\displaystyle\sum_{\ell=1}^{k}r_{\ell}\xi^{2}_{i}-f'\displaystyle\sum_{\ell=1}^{k}r_{\ell}c_0\\ 
        &=(n-1)c_{0}R\frac{1}{f'}+\vert \mathrm{Ric}\vert^2\frac{1}{f'}-f'(n-1)c_0
\end{align*}
    where we have used \eqref{excep'}, \eqref{third38'}, \eqref{eqXiRelFuncWarp''} and \eqref{scl_ric}. As a consequence,
    \begin{align}\label{eq_frst_contr}
        \vert \mathrm{Ric}\vert^2=(n-1)c_0(f')^2-(n-1)c_{0}R+f'R'.
\end{align}
On the other hand, it follows from \cite{kim3} that
	\begin{align}\label{lplacescl}
            \begin{split}
    		\Delta R&=R''-R'\displaystyle\left(\frac{(n-1)\lambda-R+\lambda_{1}}{f'}\right),
            \end{split}
	\end{align}
    and from equation $(7.1)$ of \cite[Lemma 21]{feng} it follows that
    	\begin{align}\label{lplacescl2}
		\frac{1}{2}\Delta R-\frac{1}{2}f'R'=\lambda R-|\text{Ric}|^2 
	\end{align}
    Now, once we combine \eqref{lplacescl}, \eqref{lplacescl2} and \eqref{third38'}, we obtain
\begin{align*}
    R''+R'\left(\frac{R+(n-1)c_{0}}{f'}\right)-f'R'=-2\vert \mathrm{Ric}\vert^2
\end{align*}
Proceeding, identity \eqref{frst_cnsqnc} allow to rewrite the equation above as
\begin{align}\label{eq_scnd_contr}
    2\vert \mathrm{Ric}\vert^2=4(n-1)^2c_{0}^2+2(n-1)c_{0}R+f'R',
\end{align}
where we have also combined \eqref{frst_cnsqnc} and \eqref{third38'} to get $R''=-2(n-1)^2c_{0}^2$. Putting equations \eqref{eq_frst_contr} and \eqref{eq_scnd_contr} together, a simple computation delivers that
\begin{align*}
    4(n-1)c_{0}R-2(n-1)c_{0}(f')^2-f'R'+4(n-1)^2c_{0}^{2}=0.
\end{align*}
Computing the second derivative of the equation above, and then using $R''=-2(n-1)^2c_{0}^2$, \eqref{frst_cnsqnc} and \eqref{third38'}, we conclude that $(n-1)^3c_{0}^3=0$, which can happen only if $c_{0}=0$. In this case, \eqref{frst_cnsqnc} implies $R'=0$, and inserting these last facts in \eqref{eq_frst_contr}, we obtain $\vert \mathrm{Ric}\vert^2=0$. In particular, we get $R=0$. Putting all this together, we deduce from \eqref{def_B_C} that $f'=0$, which is a contradiction.
\end{proof}

\begin{proof}[{\bf Proof of Theorem \ref{decompwarpint}}]
    From Theorem \ref{decompwarp}, $M$ is locally a multiply warped product with a one-dimensional base and $k$ fibers, counted according to Remark \ref{grmnt_wrpngfnctn}. From Lemma \ref{lem_ply_deg2_}, $\xi_{i}=h'_{i}/h_{i}$ is a root of \eqref{poly_atmst2}. Since $f$ is not constant, there are at most two distinct $\xi_{i}$. Equivalently, there are at most two fibers in the local representation given by Theorem \ref{decompwarp}.
\end{proof}

\begin{proof}[{\bf Proof of Theorem \ref{maxxnumeig-INTRO}}]
    Assume by contradiction that the Ricci tensor of $M$ has at least $4$ distinct eigenvalues in an open connected set $U\subset M$, and let $\lambda_{0},\ \lambda_{1},\ \lambda_{2}$ and $\lambda_{3}$ be three of them, pairwise distinct. Notice that $f$ is not constant in $U$. On the other hand, using Theorem \ref{decompwarp} and Theorem \ref{decompwarpint}, and shrinking $U$ if necessary, we can take it isometric to a multiply warped product with a one-dimensional base and at most two fibers. Observe that by identity \eqref{eqXiRelFuncWarp}, we must have $\xi_{1}$, $\xi_{2}$ and $\xi_{3}$, pairwise distinct, which is a contradiction. This and the density of the set of regular values of $f$ in $M$ prove the result.
\end{proof}

\section{Acknowledgments}
V. Borges was partially supported by Capes Finance Code 001, and he thanks the Mathematics Department of Universidade de Brasília, where this work was conducted. M. A. R. M. Horácio was supported by Capes Finance Code 001. J. P. dos Santos was partially supported by CNPq 313200/2025-4 and FAPDF 00193-00001678/2024-39.

\end{document}